# Dept. of Systems Engineering, ETS Montreal, Canada

### v1

@@ Abstract
A key challenge in visible-infrared person re-identification (V-I ReID) is training a backbone model capable of effectively addressing the significant discrepancies across modalities. State-of-the-art methods that generate a single intermediate bridging domain are often less effective, as this generated domain may not adequately capture sufficient common discriminant information. This paper introduces the Bidirectional Multi-step Domain Generalization (BMDG), a novel approach for unifying feature representations across diverse modalities. BMDG creates multiple virtual intermediate domains by finding and aligning body part features extracted from both I and V modalities. Indeed, BMDG aims to reduce the modality gaps in two steps. First, it aligns modalities in feature space by learning shared and modality-invariant body part prototypes from V and I images. Then, it generalizes the feature representation by applying bidirectional multi-step learning, which progressively refines feature representations in each step and incorporates more prototypes from both modalities. In particular, our method minimizes the cross-modal gap by identifying and aligning shared prototypes that capture key discriminative features across modalities, then uses multiple bridging steps based on this information to enhance the feature representation. Experiments [^1] Our code is provided in the supplementary materials and will be made public. conducted on challenging V-I ReID datasets indicate that our BMDG approach outperforms state-of-the-art part-based models or methods that generate an intermediate domain from V-I person ReID.
Keywords:
Visible-Infrared Person Re-Identification Gradual Domain Generalization Cross-Modal Image Retrieval
@@ 1 Introduction
Visible-infrared ReID (V-I ReID) is a variant of person ReID that involves matching individuals across RGB and IR cameras. V-I ReID is challenging since it requires matching individuals with significant differences in appearance between V and I modality images. In this context, V-I ReID systems must train a discriminant feature extraction backbone to encode consistent and identifiable attributes in RGB and IR images. Most state-of-the-art methods seek to learn global representations from the whole image by alignment at the image-level [38, 21] or feature-level [53, 49, 62, 33]. Fig. 1(a) illustrates the global representation approaches. Some other methods extract global modality-invariant features by disentangling them from modality-specific information [8, 48]. Global feature-based approaches cannot compare local details, resulting in the loss of important discriminant cues about the person. To address this issue and focus on the unique information in different body regions, part-based approaches (see Fig. 1(b)) extract local part-based feature representations through horizontal stripes, clustering, or attention mechanisms of spatially extracted features [52, 27, 18, 46]. However, given the significant domain discrepancies between V and I modalities and the lack of attention to these gaps at the part-level features, these methods tend to learn modality-specific attributes for each part, often leading to less discriminative feature representations.
@@
Figure 1: A comparison of training architectures for V-I ReID. Approaches based on (a) a global features representation, and (b) a local part-based representation to preserve locality and global features. (c) Approaches based on an intermediate modality generate an intermediate bridging domain. (d) Our BMGD extracts and combines prototypes from modalities in each step to gradually create multiple intermediate bridging domains.
Some recent methods leverage an intermediate modality to guide the training process such that the cross-modal domain gap is reduced [22, 54, 2, 43, 58] as shown in Fig. 1(c). These intermediate modalities may be static (e.g., gray-scale) images [54, 2], or synthesized images generated from the V or I modality [22, 43]. Although the resulting intermediate representations can leverage shared relations between I and V images and improve matching accuracy, the created images tend to contain information that is common in the background and lose discriminative ID information in the foreground. Moreover, these one-step training strategies only create one intermediate space between V and I, preventing them from controlling differences w.r.t. the primary modalities to reduce the modality gap. To effectively bridge the large cross-modal domain gap, we introduce a gradual, multi-step transformation that can preserve more ID-related information by controlling the amount of modality-specific information. Moreover, generating consecutive images results in degradation due to the loss of discriminative details accumulated over successive generations, as the images cannot provide useful information in the next steps. Therefore, our proposed method creates multiple intermediate domains in feature space by leveraging and aligning ID-informative subset of features from both modalities and then combining them.
@@
This paper introduces a novel Bidirectional Multi-step Domain Generalization (BMDG) training strategy for person V-I ReID (see Fig. 1(d)). At first, it learns to align V and I modalities and then creates multiple virtual intermediate domains between them to gradually reduce the domain gap for a backbone model in the training process. Our training strategy comprises two parts. The aligning learning module based on part-prototype [^2] We define a part prototype as a feature representation corresponding to a specific body part in a cropped person image. discovery, which extracts discriminative sub-features linked to body parts in V and I modality images. Then, the bidirectional multi-step learning progressively generates auxiliary intermediate feature representations by combining prototypes from both modalities at each step. For ID-aware and robust alignment between modalities, the extracted prototypes must meet three conditions: (1) they should be complementary with respect to other parts, (2) they should be interchangeable represent consistent parts, and (3) they should be discriminant for person matching. So, hierarchical contrastive learning is proposed to train the alignment module to extract these properties and without affecting the ReID accuracy. To leverage common semantic information between modalities and discriminant information among people, our method creates intermediate domains by mixing a growing number of body part-prototypes extracted from I and V modalities. The conditions guarantee that mixing these part-prototypes does not duplicate or degrade ID-related information of the original images.
@@
Bidirectional multi-step learning relies on auxiliary intermediate feature spaces that are created at each step, by progressively mixing semantically consistent prototypes from each modality. While prototype-based approaches such as [52, 26, 12] concentrate solely on enhancing discriminatory representation using local descriptors, BMDG aligns and mixes local part information in a gradual training process to mitigate domain discrepancies within the representation space. Unlike the Part-Mix method[20], which utilizes a fixed proportion of mixed parts for auxiliary features, our approach employs bidirectional multi-step generation for intermediate steps to bolster the robustness of V-I ReID training for large discrepancy samples. Our strategy enables effective bridging of cross-modality gaps, and allows gradual training by adjusting the proportion of exchanged information. Discriminative and contrastive regularization are used to guide the model training to detect these parts without supervision. Parts of the same individual are exchanged across modalities. Contrastive objectives are used to learn for both original and virtual features in a bidirectional way.
@@
(2) A novel hierarchical prototype learning module is introduced to align discriminative, complementary, and interchangeable part-prototypes between modalities using two-level contrastive learning. These prototypes should maintain semantic consistency to facilitate the exchange of V and I information, thus enabling the generation of intermediate bridging domains.
(3) A bidirectional multi-step learning model is proposed to progressively combine prototypes from both modalities and create intermediate feature spaces at each step. It integrates attentive prototype embedding (APE), which refines these prototypes to extract attributes that are shared between modalities, and thereby improve ReID accuracy. Indeed, BMDG can find better modality-shared attributes by optimizing the model on two levels: prototype alignment and bidirectional multi-step embedding modules.
(4) An extensive set of experiments on the challenging SYSU-MM01 [42] and RegDB [32] datasets indicate that our proposed BMDG strategy can outperform SOTA methods for V-I person ReID. They also show that our framework can be integrated into any part-based V-I ReID method and can improve the performance of cross-modal retrieval applications.
@@ 2 Related Work
(a) Global and Part-Based Representation Methods: Representation approaches focus on training a backbone model to learn a discriminant representation that captures the essential features of each modality while exploiting their shared information [50, 10, 49, 44, 25, 53]. In [53, 54, 51, 19, 33, 6], authors only use global features combined through multimodal fusion. The absence of local information limits their accuracy in more challenging cases. To improve robustness, recent methods combine global information with local part-base features [52, 27, 18, 46, 12, 20]. For example, [52, 27] divides the spatial feature map into fixed horizontal sections and applies a weighted-part aggregation. [12, 46, 45] dynamically detects regions in the spatial feature map to alleviate the misalignment of fixed horizontally divided body parts. However, these part detectors might overfit to specific parts, varying based on the modality or person. [20] address this issue using Part-Mix data augmentation to regularize the model part discovery. The bipartite graph-based correlation model between part regions has been proposed to maximize similarities [45]. While part-based methods enhance representation ability, they often overlook the shared information between similar semantic parts across different individuals or modalities. This leads to inconsistencies in the extracted part features, hindering effective comparison or exchange. Similar to the human perception system, which distinguishes individuals by comparing attributes of similar body parts (e.g., head, torso, leg, etc.) learned from a diverse set of individuals, our approach seeks to capture and leverage this shared information for accurate matching. To achieve this, our proposed BMDG method discovers different body parts as prototypes, disregarding the person’s identity, and then leverages their ID-discriminative attributes.
@@
(b) Methods Based on an Intermediate Modality: In recent years, V-I ReID methods have relied on an intermediate modality to address the significant gap between V and I modalities [11, 54, 2, 22, 43, 58, 13, 3]. Using fixed intermediate modality like grayscale [11, 54] or random channels [51, 2] can improve accuracy. [22] transform V images into a new modality to reduce the gap with I images. [43] conducted a pixel-to-pixel feature fusion operation on V and I images to build the synthetic images, and [58] used a shallow auto-encoder to generate intermediate images from both modalities. While these generated images bridge the cross-modal gap, details of the original image are lost. Accuracy is degraded in the context of multi-step learning. To address this issue, we instead extract prototypes that represent body parts and gradually use them to define multiple intermediate feature spaces.
@@
(c) Data Augmentation for Domain Generalization: Data augmentation is an effective strategy to learn domain-invariant features and improve model generalization [59]. DG methods focus on manipulating the inputs to assist in learning general representations. [60] relies on domain information to create additive noise that increases the diversity of training data while preserving its semantic information. [56, 61] uses Mixup to increase data diversity by blending examples from the different training distributions. The Representation Self-Challenging method [17] iteratively discards the dominant features from training data, aiming to improve generalization. This approach is inspired by the style transfer literature, where the feature statistics encode domain-related information. Similarly, MixStyle [61] synthesizes novel domains by mixing the feature statistics of two instances. Feature augmentation networks can transform domain styles by generating augmented examples at the feature level [53, 15, 61, 28]. [31] proposes disentangling style encodings from class categories to prevent style-biased predictions and focus more on the contents. The performance of these methods depends on whether the augmentation can help the model learn data invariance. Finally, et al. [4] relies on uniform sampling and a rejection reward, inspired by TrivialAugment [30] and TeachAugment [36], respectively.
@@
Some domain adaptation (DA) methods allow for bridging significant shifts between source and target data distributions. Gradual and multi-step (or transitive) methods [40] define the number and location of intermediate domains. [1] proposes a variant of gradual self-training that can generate pseudo-data, interpolated from the source and target domains, to define intermediate domains. Intermediate Domain Labeler [5] also creates multiple intermediate domains, where a coarse domain discriminator sorts them such that the cycle consistency of self-training is minimized. [57] proposes an auxiliary model self-training method (AuxSelfTrain) that learns models intermediate domains and gradually overcomes the shift by combining a decreasing proportion of source data while increasing the proportion of target data. In this paper, we iteratively generate features to define multiple intermediate virtual domains by gradually mixing various body part prototypes extracted from I and V modalities. This DG approach gradually reduces the cross-modality gap over multiple intermediate training steps.
@@ 3 Proposed Method
Figure 2: (a) Overall BMDG training architecture comprises two parts. The prototype learning module (left) extracts body part prototype representations from V and I images. The bidirectional multi-step learning module (right) extracts discriminant features using multiple intermediate domains created by mixing prototype information. (b) Prototype discovery (PD) architecture mines prototypes from spatial features, and Hierarchical contrastive learning (HCL) encourages the prototypes to focus on similar semantics for all individuals without losing ID-discriminative information.
Preliminaries: For the training process, let us assume a multimodal dataset for V-I ReID, represented by $\mathcal{V}=\{x^{j}_{v},y_{v}^{j}\}_{j=1}^{N_{v}}$ and =$\mathcal{I}=\{x^{j}_{i},y_{i}^{j}\}_{j=1}^{N_{i}}$ which contains sets of V and I images from $C_{y}$ distinct individuals, with their ID labels. V-I ReID systems seek to match cropped person images captured from V and I cameras by training a deep backbone model to encode modality-invariant person embedding, denoted by $\mathbf{f}_{v}$ and $\mathbf{f}_{i}$. Given images of the same person, the objective is to maximize similarity:
$\max S(\mathbf{f}_{v}^{n},\mathbf{f}_{i}^{l})\;\forall n\in\{1,\dots,N_{v}\}\text{ and }l\in\{1,\dots,N_{i}\}\text{ with }y_{v}^{n}=y_{i}^{l},\vspace{-0.3cm}$ (1)
and for different people, we seek to minimize similarity:
$\min S(\mathbf{f}_{v}^{n},\mathbf{f}_{i}^{l})\;\forall n\in\{1,\dots,N_{v}\}\text{ and }l\in\{1,\dots,N_{i}\}\text{ with }y_{v}^{n}\neq y_{i}^{l}\vspace{-0.1cm}$ (2)
@@ 3.1 Prototype Learning Module
To align pedestrians’ sub-features between different images in I and V modalities, this module is proposed to discovery different body-part attributes from diverse regions and then learns to align these prototypes by making them complementary, interchangeable, and discriminative, as discussed in Section 1.
(a) Prototype Discovery (PD): This module is proposed to extract the prototypes and discover several discriminative regions on the backbone feature maps. At first, the PD predicts a pixel-wise masking score (probability) $M\in\mathbb{R}^{H\cdot W\times K}$ for each prototype from the features map $\mathbf{F}_{m}\in\mathbb{R}^{H\cdot W\times d}$ where $H,W,d$ are the feature sizes, and $K$ is the total number of prototypes. $m$ is modality index which indicates $v$ as visible or $i$ as infrared domain. Then, it computes each prototype feature $\mathbf{p}_{m}^{k}$ by weighted aggregating of features in all pixels as:
@@
where for each pixel $u\in U=\{1,\dots,H\cdot W\}$, we have $\sum_{k=1}^{K}[M_{m}^{k}]_{u}=1$. This module (illustrated in Fig. 2(b) uses a shallow U-Net[34] architecture to create a region mask for each prototype. In the down-sample block, a convolution receptive field is applied on masks to make them lose locality by leveraging their neighbor pixels. At the same time, the up-sample propagates these high-level, coarsely localized masks into each original size. Skip-connection is used to re-inject the local details lost in the down-sample phase.
(b) Hierarchical Contrastive Learning: To ensure the part prototypes represent the diverse attributes of each person, they must be non-redundant, and each part must be independent of the others. This independence can be formulated that the mutual information ($MI$) between the distribution of random variables describing parts as $P^{k}$ should be zero as:
$MI(P^{k},P^{q})=0\;\forall k,q\in\{1,\dots,K\},\;k\neq q,$ (4)
@@
$\min S(\mathbf{p}^{k},\mathbf{p}^{q}))\;\forall k,q\in\{1,\dots,K\},\;k\neq q,$ (5)
@@
where $\mathbf{p}^{k}$ and $\mathbf{p}^{q}$ are features for part $k$ and $q$, respectively. To assure that the prototype features are semantically consistent around body parts can be aligned in different samples, we maximize the similarity of $\mathbf{p}^{k}$ and all other prototypes with the same index in the training batch as $\mathbf{\hat{p}}^{k}$.
@@
To select samples in each batch with a positive prototype index, there are two options: the same parts belonging to the same or different persons (inter-person) as $\Psi^{k}$, and the same parts belonging to the same person with identity $y$ (intra-person) as $\Gamma^{k}_{y}$ ( see Fig. 2.b). In a deep ReID model, the goal is to extract dissimilar features in the representation space for two different persons, so pulling apart the same prototypes from these two persons may disrupt this objective. It is preferable to let them encode ID-related attributes. In contrast, the early backbone layers focus on semantically similar prototype attributes, disregarding a person’s identity. Therefore, two contrastive losses as $\mathcal{L}_{\text{lfc}}$ and $\mathcal{L}_{\text{hfc}}$ are employed to provide inter-peeson information at a low-level and intra-person at a high-level:
$\mathcal{L}_{\text{lfc}}=-\sum_{k=1}^{K}\sum_{\mathbf{\hat{p}^{\prime}}\in\Psi^{k}}\log\frac{e^{\mathbf{p}^{\prime k}\cdot\mathbf{\hat{p}^{\prime}}/\tau}}{e^{\mathbf{p}^{\prime k}\cdot\mathbf{\hat{p}^{\prime}}/\tau}+\sum_{q\neq k}e^{\mathbf{p^{\prime}}^{k}\cdot\mathbf{p^{\prime}}^{q}/\tau}},$ (6)
$\mathcal{L}_{\text{hfc}}=-\sum_{y=1}^{C_{y}}\sum_{k=1}^{K}\sum_{\mathbf{\hat{p}}\in\Gamma^{k}_{y}}\log\frac{e^{\mathbf{p}^{k}\cdot\mathbf{\hat{p}}/\tau}}{e^{\mathbf{p}^{k}\cdot\mathbf{\hat{p}}/\tau}+\sum_{q\neq k}e^{\mathbf{p}^{k}\cdot\mathbf{p}^{q}/\tau}},$ (7)
@@
where $\mathbf{p}_{m}^{\prime k}$ are low-level part prototype features computed same as Eq. (3) but extracted from lower-level layers of the feature backbone. Unlike contrastive learning introduced in [24, 9] that are used only for final features, the HCL optimizes in two levels as low-level and high-level features to force the model to cluster inputs based on part and then cluster inside identities, as illustrated in Fig. 2(b). For example, $\mathcal{L}_{\text{lfc}}$ encourages the low-level prototype of legs for two persons be similar and differ from other prototypes such as torso. While $\mathcal{L}_{\text{hfc}}$ encourages the high-level legs prototypes only differs from other prototypes in that person. Also, minimizing the distance between features of pixels belonging to the same part occurrence encourages the model to extract more similar features for each prototype:
$\mathcal{L}_{\text{part}}=\sum_{k=1}^{K}\sum_{u\in U}[M_{m}^{k}\|\mathbf{p}_{m}^{k}-\mathbf{F}_{m}\|]_{u}.\vspace{-0.15cm}$ (8)

### v2

# Bidirectional Multi-Step Domain Generalization for Visible-Infrared Person Re-Identification

@@ Abstract
A key challenge in visible-infrared person re-identification (V-I ReID) is training a backbone model capable of effectively addressing the significant discrepancies across modalities. State-of-the-art methods that generate a single intermediate bridging domain are often less effective, as this generated domain may not adequately capture sufficient common discriminant information. This paper introduces Bidirectional Multi-step Domain Generalization (BMDG), a novel approach for unifying feature representations across diverse modalities. BMDG creates multiple virtual intermediate domains by learning and aligning body part features extracted from both I and V modalities. In particular, our method aims to minimize the cross-modal gap in two steps. First, BMDG aligns modalities in the feature space by learning shared and modality-invariant body part prototypes from V and I images. Then, it generalizes the feature representation by applying bidirectional multi-step learning, which progressively refines feature representations in each step and incorporates more prototypes from both modalities. Based on these prototypes, multiple bridging steps enhance the feature representation. Experiments [^1] Our code is available at: alehdaghi.github.io/BMDG conducted on V-I ReID datasets indicate that our BMDG approach can outperform state-of-the-art part-based and intermediate generation methods, and can be integrated into other part-based methods to enhance their V-I ReID performance.
@@ 1 Introduction
V-I ReID is a variant of person ReID that involves matching individuals across RGB and IR cameras. V-I ReID is challenging since it requires matching individuals with significant differences in appearance between V and I modality images. In this context, V-I ReID systems must train a discriminant feature extraction backbone to encode consistent and identifiable attributes in RGB and IR images. Most state-of-the-art methods seek to learn global representations from the whole image by alignment at the image-level [32, 19] or feature-level [47, 44, 59, 28] (see Fig. 1(a)). Other methods extract global modality-invariant features by disentangling them from modality-specific information [7, 42]. However, global feature-based approaches cannot compare local discriminative attributes, resulting in the loss of important discriminant cues about the person. To address this issue and focus on the unique information in different body regions, part-based approaches (see Fig. 1(b)) extract local fine-grained feature representations through horizontal stripes, clustering, or attention mechanisms of spatially extracted features [46, 25, 17, 40]. However, given the domain discrepancies between the V and I modalities and the lack of matching cross-modal parts during training, these methods often learn modality-specific attributes for each part. This typically results in feature representations that are less modality-invariant, and thus provide limited effectiveness for cross-modal matching tasks.
@@
Figure 1: A comparison of training architectures for V-I ReID. Approaches based on (a) a global features representation, and (b) a local part-based representation to preserve locality and global features. (c) Generation of an intermediate bridging domain to guide training. (d) Our BMGD extracts and combines prototypes from modalities in each step to gradually create multiple intermediate bridging domains.
Some methods leverage an intermediate modality to guide the training process such that the cross-modal domain gap is reduced [20, 48, 2, 37, 54] (see Fig. 1(c)). Intermediate modalities may be static images (e.g., gray-scale) [48, 2] or created images from a generative module learned by I and V modalities [20, 37, 54]. The resulting intermediate representations can leverage shared relations between I and V images and improve robustness to modality changes. However, they generate images that tend to contain shared non-discriminative information, resulting in the loss of ID-aware information in the intermediate domain. Moreover, one-step training strategies only create one intermediate space between V and I. This cannot provide sufficient bridging information when modalities have significant discrepancies. Single-step approaches cannot effectively deal with these discrepancies during training. Therefore, we advocate for linking a series of intermediate domains with progressively smaller gaps.
To effectively bridge the large cross-modal domain gap while preserving local ID-discriminative information, we introduce Bidirectional Multi-step Domain Generalization (BMDG), a gradual, multi-step, part-based training strategy for person V-I ReID (see Fig. 1(d)). At first, this strategy identifies similar ID-aware part prototypes [^2] We define a part prototype as a feature representation corresponding to a specific body part in a cropped image of a person. for each modality, and then it progressively creates intermediate steps by controlling the number of the part prototypes to be mixed. Generated images can become degraded, as discriminative details are lost over consecutive generation steps, making these images less useful for training. Therefore, our proposed BMDG creates multiple intermediate domains in the feature space by aligning, and then combining ID-informative subsets of features from both modalities.
@@
Our BMGD training strategy is comprised of two parts. The prototype alignment module learns to identify and align part-prototype representations that contain discriminative information linked to body parts in I and V images. Then, the bidirectional multi-step learning progressively extracts auxiliary intermediate feature representations by mixing prototypes from both modalities at each step. For ID-aware and robust alignment between modalities, the extracted part-prototypes must meet three conditions: (1) they should be complementary to other parts, (2) they should be interchangeable and represent consistent parts, and (3) they should be discriminant for person matching. Hierarchical contrastive learning is proposed to train the prototype alignment module such that these properties are respected without affecting ReID accuracy. Our novel idea is to progressively create intermediate bridging domains by mixing a growing number of body part prototypes extracted from I and V modalities. This allows leveraging common semantic information between modalities, along with discriminative information among individuals.
@@
Bidirectional multi-step learning relies on auxiliary intermediate feature spaces that are created at each step, by progressively mixing semantically consistent prototypes from each modality. While prototype-based approaches [46, 24, 12, 58] concentrate solely on enhancing discriminatory representation using local descriptors, BMDG aligns and mixes local part information in a gradual training process to mitigate domain discrepancies within the representation space. Unlike the Part-Mix method [18], which utilizes a fixed proportion of mixed parts for auxiliary features, our approach employs bidirectional multi-step generation for multiple intermediate steps. This bolsters the robustness of V-I ReID training for samples with large discrepancies, especially at the beginning of training when the parts are inconsistent. Our strategy enables effective bridging of cross-modality gaps and allows gradual training by adjusting the proportion of information to be exchanged. We also propose a hierarchical prototype learning strategy, which is crucial to defining meaningful bridging steps.
@@
(2) A novel hierarchical prototype learning module is introduced to align discriminative and interchangeable part prototypes between modalities using two-level contrastive learning. These prototypes maintain semantic consistency to facilitate the exchange of V and I information, thus enabling the generation of intermediate bridging domains.
(3) A bidirectional multi-step learning model is proposed to progressively combine prototypes from both modalities and create intermediate feature spaces at each step. It integrates attentive prototype embedding (APE), which refines these prototypes to extract attributes that are shared between modalities, thereby improving ReID accuracy. Indeed, BMDG can find better modality-shared attributes by optimizing the model on two levels: prototype alignment and bidirectional multi-step embedding modules.
(4) Extensive experiments on the challenging SYSU-MM01 [36], RegDB [27] and LLCM[52] datasets indicate that BMDG outperforms state-of-the-art methods for V-I person ReID. They also show that our framework can be integrated into any part-based V-I ReID method and can improve the performance of cross-modal retrieval applications.
@@ 2 Related Work
(a) Global and Part-Based Representation Methods: Representation approaches focus on training a backbone model to learn a discriminant representation that captures the essential features of each modality while exploiting their shared information [43, 9, 44, 38, 23, 47]. In [47, 48, 45, 28, 5], authors only use global features through multimodal fusion. The absence of local information limits their accuracy in more challenging cases. To improve robustness, recent methods combine global information with local part-base features [46, 25, 40, 12, 18]. For example, [46, 25] divides the spatial feature map into fixed horizontal sections and applies a weighted-part aggregation. [12, 40, 39, 58] dynamically detects regions in the spatial feature map to alleviate the misalignment of fixed horizontally divided body parts.
However, these part detectors might over-focus on specific parts, varying based on the modality or person. [18] addresses this issue using Part-Mix data augmentation to regularize model part discovery. Additionally, parts may not correspond to semantically similar regions across different images. To tackle this, a bipartite graph-based correlation model between part regions has been proposed to maximize similarities [39]. While part-based methods enhance representation ability, they often overlook shared information between similar semantic parts across different individuals or modalities, leading to inconsistencies in extracted part features. This hinders effective comparison or exchange. Similar to the human perception system, distinguishing individuals by comparing attributes of similar body parts (e.g., head, torso, leg) learned from a diverse set of individuals, our approach seeks to capture and leverage this shared information for accurate matching. To achieve this, BMDG discovers different body parts as prototypes, disregarding the person’s identity using hierarchical contrastive learning, and then leverages their ID-discriminative attributes for each part through part-based cross-entropy loss.
@@
(b) Methods Based on an Intermediate Modality: In recent years, V-I ReID methods have relied on an intermediate modality to address the significant gap between V and I modalities [11, 48, 2, 20, 37, 54, 13, 3]. Using fixed intermediate modality like grayscale [11, 48] or random channels [45, 2] can improve accuracy. [20] transform V images into a new modality to reduce the gap with I images. [37] conducted a pixel-to-pixel feature fusion operation on V and I images to build the synthetic images, and [54] used a shallow auto-encoder to generate intermediate images from both modalities. While these generated images bridge the cross-modal gap, details of the image are lost. Accuracy is degraded in the context of multi-step learning. To address this, we extract body part prototypes and gradually use them to define multiple intermediate feature spaces.
@@
(c) Domain Generalization and Adaptation: Data augmentation is an effective strategy to learn domain-invariant features and improve model generalization [55]. DG methods focus on manipulating the inputs to assist in learning general representations. [56] relies on domain information to create additive noise that increases the diversity of training data while preserving its semantic information. [50, 57] uses Mixup to increase data diversity by blending examples from the different training distributions.
@@
Some domain adaptation (DA) methods allow for bridging significant shifts between source and target data distributions. Gradual and multi-step (or transitive) methods [34] define the number and location of intermediate domains. Intermediate Domain Labeler [4] also creates multiple intermediate domains, where a coarse domain discriminator sorts them such that the cycle consistency of self-training is minimized. IDM[10] mixes entire latent features to create an intermediate domain that lies on the shortest path between the target and source. In fact, by utilizing appropriate intermediate domains, the source knowledge can be better transferred to the target domain. By mixing the entire hidden features, this domain may lose the local ID-discriminative information and reduce the diversity of part features. This paper reduces the cross-modality gap by exchanging features between domains and creating intermediate virtual domains through gradual mixing during training.
@@ 3 Proposed Method
Figure 2: (a) Overall BMDG training architecture comprises two parts. The prototype alignment module (left) extracts body part prototype representations from V and I images. The bidirectional multi-step learning module (right) extracts discriminant features using multiple intermediate domains created by mixing prototype information. (b) Prototype discovery (PD) architecture mines prototypes from spatial features, and Hierarchical contrastive learning (HCL) encourages the prototypes to focus on similar semantics for all individuals without losing ID-discriminative information.
Preliminaries: For the training process, let us assume a multimodal dataset for V-I ReID, represented by $\mathcal{V}=\{x^{j}_{v},y_{v}^{j}\}_{j=1}^{N_{v}}$ and =$\mathcal{I}=\{x^{j}_{i},y_{i}^{j}\}_{j=1}^{N_{i}}$ which contains sets of V and I images from $C_{y}$ distinct individuals, with their ID labels. V-I ReID systems seek to match images captured from V and I cameras by training a deep backbone model to encode modality-invariant person embedding, denoted by $\mathbf{f}_{v}$ and $\mathbf{f}_{i}$. Given V and I images, the objective is to:
[…]
@@ 3.1 Prototype Learning Module
To align persons’ sub-features between different images in I and V, this module is used to discover different body-part attributes from diverse regions and learn to align these prototypes by making them complementary, interchangeable, and discriminative, as discussed in Section 1.
$\mathcal{H}\colon$ $\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{K\times d}$ (2)
$x\sim X\mapsto\mathcal{H}(x)\sim P$ (3)
(a) Prototype Discovery (PD). This module is proposed to extract the prototypes and discover several discriminative regions on the backbone feature maps. At first, the PD predicts a pixel-wise masking score (probability) $M\in\mathbb{R}^{H\cdot W\times K}$ for each prototype from the features map $\mathbf{F}_{m}\in\mathbb{R}^{H\cdot W\times d}$ where $H,W,d$ are the feature sizes, and $K$ is the total number of prototypes. $m$ is modality index which indicates $v$ as visible or $i$ as infrared domain. Then, it computes each prototype vector $\mathbf{p}_{m}^{k}$ by weighted aggregating of features in all pixels as:
@@
where for each pixel $u\in U=\{1,\dots,H\cdot W\}$, we have $\sum_{k=1}^{K}[M_{m}^{k}]_{u}=1$. This module (illustrated in Fig. 2(b) uses a shallow U-Net[29] architecture to create a region mask for each prototype. In the down-sample block, a convolution receptive field is applied on masks to make them lose locality by leveraging their neighbor pixels. At the same time, the up-sample propagates these high-level, coarsely localized masks into each original size. Skip connections are used to re-inject the local details lost in the down-sample phase.
(b) Hierarchical Contrastive Learning: To ensure the part prototypes represent the diverse attributes of each person, they must be non-redundant, and be independent of the others. This independence can be formulated that the mutual information ($MI$) between the distribution of random variables describing parts as $P^{k}$ should be zero as:
$\text{MI}(P^{k};P^{q})=0\;\forall k,q\in\{1,\dots,K\},\;k\neq q,\vspace{-0.15cm}$ (5)
@@
$\min S\{(\mathbf{p}^{k},\mathbf{p}^{q})\}\;\forall k,q\in\{1,\dots,K\},\;k\neq q,\vspace{-0.15cm}$ (6)
@@
where $\mathbf{p}^{k}$ and $\mathbf{p}^{q}$ are features for part $k$ and $q$, respectively. To ensure that the prototype features are semantically consistent around body parts that can be aligned in different samples, we maximize the similarity of $\mathbf{p}^{k}$ and all other prototypes with the same index in the training batch as $\mathbf{\hat{p}}^{k}$.
@@
To select samples in each batch with a positive prototype index, there are two options: the same parts belonging to the same or different persons (inter-person) as $\Psi^{k}$, and the same parts belonging to the same person with identity $y$ (intra-person) as $\Gamma^{k}_{y}$ ( see Fig. 2.b). Since the goal is to extract dissimilar features in the representation space for two different persons, pulling apart the same prototypes from these two persons may disrupt this objective. Therefore, two contrastive losses as $\mathcal{L}_{\text{lc}}$ and $\mathcal{L}_{\text{hc}}$ are employed to provide inter-person information at a low-level and intra-person at a high-level:
$\mathcal{L}_{\text{lc}}=-\sum_{k=1}^{K}\sum_{\mathbf{\hat{p}^{\prime}}\in\Psi^{k}}\log\frac{e^{\mathbf{p}^{\prime k}\cdot\mathbf{\hat{p}^{\prime}}/\tau}}{e^{\mathbf{p}^{\prime k}\cdot\mathbf{\hat{p}^{\prime}}/\tau}+\sum_{q\neq k}e^{\mathbf{p^{\prime}}^{k}\cdot\mathbf{p^{\prime}}^{q}/\tau}},$ (7)
$\mathcal{L}_{\text{hc}}=-\sum_{y=1}^{C_{y}}\sum_{k=1}^{K}\sum_{\mathbf{\hat{p}}\in\Gamma^{k}_{y}}\log\frac{e^{\mathbf{p}^{k}\cdot\mathbf{\hat{p}}/\tau}}{e^{\mathbf{p}^{k}\cdot\mathbf{\hat{p}}/\tau}+\sum_{q\neq k}e^{\mathbf{p}^{k}\cdot\mathbf{p}^{q}/\tau}},\vspace{-0.15cm}$ (8)
@@
where $\mathbf{p}_{m}^{\prime k}$ are low-level part prototype features vector computed same as Eq. (4) but extracted from lower-level layers of the feature backbone. Unlike contrastive learning introduced in [22, 8] that are used only for final features, the HCL optimizes at two levels as low-level and high-level features to force the model to cluster inputs based on parts and then cluster inside identities, as illustrated in Fig. 2(b). For example, $\mathcal{L}_{\text{lc}}$ encourages the low-level prototype of legs for two persons to be similar and differ from other prototypes such as the torso. While $\mathcal{L}_{\text{hc}}$ encourages the high-level legs prototypes only differ from other prototypes in that person. Also, minimizing the distance between features of pixels belonging to the same part occurrence encourages the model to extract more similar features for each prototype:
$\vspace{-0.15cm}\mathcal{L}_{\text{c}}=\sum_{k=1}^{K}\sum_{u\in U}[M_{m}^{k}\|\mathbf{p}_{m}^{k}-\mathbf{F}_{m}\|]_{u}.\vspace{-0.15cm}$ (9)